If line 3x + ky = 9, passes through the point (1,-2), find k.
Answer by jim_thompson5910(35256)

Plug in (x,y) = (1,-2) and solve for k

3x + ky = 9
3(1) + k(-2) = 9
3 - 2k = 9
-2k = 9-3
-2k = 6
k = 6/(-2)
k = -3

So the answer is k = -3